OpenStudy (anonymous):

Solve for x. State any restrictions on the variables. a(x-3)+8=b(x-1)

OpenStudy (dls):

a(x-3)+8=b(x-1) a(x-3)+8-b(x-1)=0 ax-3a+8-bx+b=0 x(a-b)-3a+b+8=0 \[\Huge x=\frac{3a-b-8}{a-b}\] Restrictions on variable : \[\Huge a \neq b\] otherwise function becomes indeterminate.